OpenStudy (anonymous):

A number being successively divided by 3, 5 and 8 leaves remainders 1, 4 and 7 respectively. Find the remainders if the order of divisors be reversed.

OpenStudy (anonymous):

6,4,2

OpenStudy (loser66):

I don't understand the question. what does "the order of divisors be reversed" mean?

OpenStudy (loser66):

Let the number is n, then n = 1 [mod3] n =4 [mod5] n= 7 [mod 8] then? what is next? @mathmale

OpenStudy (anonymous):

A number when successively divided by 3,5,8 leaving remainders 1,4,7 can be written as 3(5(8k+7)+4)+1 =>120k+118 when this number is dived by 8 i.e (120k+118)/8=15k+14 remainder 6 (15k+14)/5=3k+2 remainder 4 (3k+2)/3=k remainder 2 And D) 6,4,2

OpenStudy (loser66):

got it, thank you